# PyraTrans: Attention-Enriched Pyramid Transformer for Malicious URL Detection

### v2

@@ IV-C Multi-scale Learning
IV-D Spatial Pyramid Attention
@@
TABLE IV: Detection results of our proposed method vary with the number of layers employed
# layers(count) Accuracy Precision Recall F1-score AUC
@@ V Experiments
Figure 5: Detection results of selected baseline methods and our proposed method on GramBeddings dataset.
@@
V-A Evaluate the Effectiveness of Multi-level Features
@@
We create a training corpus from the GramBeddings dataset, randomly selecting 128k URLs from the 80k available in the training set while maintaining the proportion of malicious and benign URLs. We also randomly sample 32k URLs for testing and validation respectively. The performance of various stacked configurations and the incremental gains are shown in Table IV. We observe a consistent improvement in evaluation metrics as we incrementally incorporate additional layers, namely the last 2, 3, 4, and 5 layers. Ultimately, the stacking of 12 embedding output layers achieves the best performance in URL detection tasks, demonstrating the effectiveness of integrating both lower and deeper layers in the model architecture.
V-B Comparison with Baselines
In this section, we compare the performance of PyraTrans with baselines using binary and multi-class malicious URL detection.
TABLE V: Detection results of selected baseline methods and our proposed method on Mendeley dataset.
@@ V-B Comparison with Baselines
Our 0.9834 0.7632 0.3717 0.5000 0.8550
TABLE VI: Cross-dataset performance generalization
@@ V-B1 Binary classification
Results on Grambeddings dataset: As shown in Figure 5, our proposed method achieves superior performance over the baseline method on the Grambeddings dataset, regardless of the size of the training set. Remarkably, our model demonstrates high proficiency even with scarce training data.
@@
As shown in Table V, our method exhibits significant advantages in class-imbalanced scenarios compared to other approaches. With just 1% of training data, our model achieves an accuracy of 0.9837, with a Precision of 0.7632, surpassing the best baseline Precision of 0.1808. The F1 score of our method is four times higher than the best baseline performance. As training data increases, our F1 score reaches 0.7027, substantially exceeding the baseline peak of 0.5515, with a Recall 15.3 higher than the best baseline result. These experiments demonstrate the substantial improvements our method brings in identifying malicious URL samples, significantly reducing false positives. Moreover, the high AUC reflects our model’s confidence in the identified samples, indicating its accurate capture of key differences between malicious and benign samples.
@@
We observe that while URLNet achieves high accuracy on larger datasets, it suffers from a high false negative rate, particularly on smaller datasets. Notably, URLNet fails to identify any malicious URLs when the training data size is reduced to below 5%. In contrast, Grambeddings shows significant sensitivity to data, with its performance in detecting malicious URLs varying greatly across different training sample sizes. For instance, at 60% training data, its Precision drops to 0.6137, and at 40%, its Recall falls to just 0.0817. Compared to these, our method demonstrates consistent and reliable performance across various data scales, indicating its robustness to small-scale and class-imbalanced data scenarios.
V-B2 Multi-classification
@@
To evaluate our method in the context of complex cyber threats, we conduct a multi-class classification experiment, using a Kaggle 2 dataset [36] with four URL categories: benign (428,079), defacement (95,306), phishing (94,086), and malicious (23,645). Fig. 6 shows the results of our model and the baseline methods.
@@
Fig.6 illustrates the performance of our method and the baseline methods across all four categories. Our method surpasses the baseline methods in each metric. The average ROC curve highlights our method’s efficacy with a TPR of almost 90% at a low FPR of 0.001, surpassing other methods that achieve around 75%. Grambeddings struggled in recognizing negative samples from various categories, resulting in a 50% F1 score for defacement and phishing URLs and an overall accuracy of 83.91%. URLNet achieved an overall accuracy of 97.07%, falling short of our model’s 98.57%. These results demonstrate the robustness and effectiveness of our method in complex multi-class classification tasks, indicating its potential as a promising solution for malicious URL detection in cybersecurity.
V-C Cross-dataset Testing
@@
The results, as shown in Table VI, indicate a marked decline in performance of baseline methods on data not included in their training set, with URLNet’s accuracy dropping to 0.8823 and Grambeddings’ accuracy reducing to 0.5214 and 0.4647, respectively. In contrast, our model maintained high accuracy on external datasets, achieving an AUC of 0.9705. This underscores that the knowledge gained from one dataset by our method can be effectively generalized to others, even if the data was collected much later than our model’s pretraining period. This demonstrates the adaptability and long-term applicability of our approach to different data environments.
Figure 6: Detection results of selected baseline methods and our proposed method on multiple classification dataset.
V-D Evaluation against Adversarial Attacks
@@
As illustrated in Figure 7 and Table VII, baseline methods experience a significant decline in performance under adversarial sample attacks. The accuracy of URLNet decreases to 76.10%, and that of Grambeddings to 70.18%, while our model maintains an accuracy above 90%, with an AUC of 98.39%, exceeding URLNet by about 20%. At a fixed FPR of 0.01, our method achieves a TPR of nearly 75%, more than double the TPR of baseline methods, both under 30%. These results indicate the robustness of our approach to adversarial sample attacks, suggesting increased effectiveness in preventing malicious attack evasion in real-world scenarios.
Figure 7: Area under ROC curve under adversarial attack.
TABLE VII: Performance under adversarial attack.
@@
Our 0.9023 0.9738 0.8104 0.8846 0.9839
V-E Case Study
We conduct a series of case studies applying our detection model to active malicious web pages to evaluate its practical utility. In November 2023, we crawled 30 active phishing URLs reported and verified on PhishTank and tested them with our model trained on 30% of the Grambeddings dataset. For comparison, tests were also conducted with the best-performing URLNet and Grambeddings models, trained on the same dataset. Results indicate that our model detected all malicious URLs with 100% accuracy, while Grambeddings misclassified 7 URLs, resulting in 76% accuracy, and URLNet misclassified 4 URLs, with an accuracy of 86%. Table VIII lists URLs that were incorrectly classified, in order to provide a detailed perspective on the performance of each model.
VI Discussion
TABLE VIII: Cross-dataset performance generalization
@@ VI Discussion
https://bafybeibfyqcvrjmwlpipqkdyt2xr46cea7ldciglcbybfwtk7cieugcj3e.ipfs.infura-ipfs.io
@@
    End-to-End Architecture: ur proposed method is an end-to-end network utilizing pretrained CharBERT, which requires no manual feature initialization. It directly processes raw URLs and generates character-aware subword token embeddings. In contrast, previous studies typically necessitated manual initialization of character and word-level representations and relied on dual-path neural networks. Our approach streamlines the processing workflow while maintaining efficient feature extraction capabilities.
@@
    Evaluation Metrics: Our experiments show that the proposed method significantly improves accuracy, robustness, and generalizability, consistently delivering stable and effective detection across various testing scenarios. Meanwhile, leading baseline methods, such as URLNet and Gramembedings, displayed clear weaknesses: URLNet struggled in class-imbalanced scenarios with small datasets, while Gramembedings faced significant performance fluctuation. Furthermore, these methods varied in their effectiveness in generalization and adversarial robustness tests. This highlights the need for a comprehensive performance evaluation system that goes beyond specific experimental setups.
@@
VII Conclusion
We have proposed a novel transformer-based and pyramid feature learning system called PyraTrans for malicious URL detection. Our method effectively leverages knowledge transfer from pretrained models to URL contexts, dynamically integrates character and subword-level features, and incorporates three closely integrated feature learning modules for URL feature extraction. The key contributions of our approach are: 1) enabling end-to-end learning from raw URL strings without manual preprocessing; 2) adopting an interactive subword and character-level feature learning network architecture for improved character-aware subword representations; 3) conducting effective multi-level and multi-scale URL feature learning based on our proposed lightweight feature learning modules, addressing inherent limitations of the Transformer in local feature extraction and spatial awareness. We conduct extensive experiments on various URL datasets, demonstrating that our method consistently outperforms existing state-of-the-art baseline methods and produces stable decisions across scenarios. Furthermore, our method exhibits superior generalization and robustness in cross-dataset detection and adversarial sample attacks, enhancing its reliability in practical applications. We also provide a case study with comparative analysis to demonstrate the practical value of our method.

### v3

@@ 4.3 Multi-scale Learning
4.4 Spatial Pyramid Attention
@@
Table 4: Detection results of TransURL vary with the number of layers employed.
 layers(count) Accuracy Precision Recall F1-score AUC
@@ 5 Experiments
Figure 5: Detection results of baseline methods and TransURL on GramBeddings dataset.
@@
5.1 Evaluate the Effectiveness of Multi-level Features
@@
We create a training corpus from the GramBeddings dataset, randomly selecting 128k URLs from the 80k available in the training set while maintaining the proportion of malicious and benign URLs. We also randomly sample 32k URLs for testing and validation respectively. The performance of various stacked configurations and the incremental gains are shown in Table 4. We observe a consistent improvement in evaluation metrics as we incrementally incorporate additional layers, namely the last 2, 3, 4, and 5 layers. Ultimately, the stacking of 12 embedding output layers achieves the best performance in URL detection tasks, demonstrating the effectiveness of integrating both lower and deeper layers in the model architecture.
5.2 Comparison with Baselines
In this section, we compare the performance of TransURL with baselines using binary and multi-class malicious URL detection.
Table 5: Detection results of baseline methods and TransURL on Mendeley dataset.
@@ 5.2 Comparison with Baselines
TransURL 0.9834 0.7632 0.3717 0.5000 0.8550
Table 6: Cross-dataset performance generalization.
@@ 5.2.1 Binary classification
Results on GramBeddings dataset: As shown in Fig. 5, our proposed method achieves superior performance over the baseline method on the GramBeddings dataset, regardless of the size of the training set. Remarkably, our model demonstrates high proficiency even with scarce training data.
@@
As shown in Table 5, TransURL exhibits significant advantages in class-imbalanced scenarios compared to other approaches. With just 1% of training data, TransURL achieves an accuracy of 0.9837, with a Precision of 0.7632, surpassing the best baseline Precision of 0.1808. The F1 score of TransURL is four times higher than the best baseline performance. As training data increases, our F1 score reaches 0.7027, substantially exceeding the baseline peak of 0.5515, with a Recall 15.3 higher than the best baseline result. These experiments demonstrate the substantial improvements TransURL brings in identifying malicious URL samples, significantly reducing false positives. Moreover, the high AUC (Area Under the Curve) reflects our model’s confidence in the identified samples, indicating its accurate capture of key differences between malicious and benign samples.
@@
We observe that while URLNet achieves high accuracy on larger datasets, it suffers from a high false negative rate, particularly on smaller datasets. Notably, URLNet fails to identify any malicious URLs when the training data size is reduced to below 5%. In contrast, GramBeddings shows significant sensitivity to data, with its performance in detecting malicious URLs varying greatly across different training sample sizes. For instance, at 60% training data, its Precision drops to 0.6137, and at 40%, its Recall falls to just 0.0817. Compared to these, TransURL demonstrates consistent and reliable performance across various data scales, indicating its robustness to small-scale and class-imbalanced data scenarios.
These notable performance improvement with limited training data can be attributed to several key factors. Firstly, TransURL utilizes multi-layer transformer encoding, enabling efficient capture of long-range dependencies and intricate patterns within URLs. This capability is particularly advantageous in scenarios with scarce training data, as the model demonstrates superior generalization from fewer examples. Secondly, by integrating multi-scale pyramid features, TransURL is capable of analyzing URLs at various granularities. This multi-scale approach ensures the detection of critical features at different levels, thereby enhancing the model’s ability to distinguish between benign and malicious URLs even with limited training data. Moreover, we employ transfer learning techniques, wherein TransURL is pre-trained on a larger textual dataset before being fine-tuned for the specific task of URL classification. This pre-training phase equips the model with a robust foundational understanding, significantly enhancing its performance during fine-tuning with limited data.
5.2.2 Multi-classification:
@@
To evaluate TransURL in the context of complex cyber threats, we conduct a multi-class classification experiment, using a Kaggle 2 dataset [30] with four URL categories: benign (428,079), defacement (95,306), phishing (94,086), and malicious (23,645). Fig. 6 shows the results of our model and the baseline methods.
@@
Fig.6 illustrates the performance of our method and the baseline methods across all four categories. TransURL surpasses the baseline methods in each metric. The average ROC (Receiver Operating Characteristic curve) curve highlights our method’s efficacy with a TPR (True Positive Rate) of almost 90% at a low FPR(False Positive Rate) of 0.001, surpassing other methods that achieve around 75%. GramBeddings struggled in recognizing negative samples from various categories, resulting in a 50% F1 score for defacement and phishing URLs and an overall accuracy of 83.91%. URLNet achieved an overall accuracy of 97.07%, falling short of our model’s 98.57%. These results demonstrate the robustness and effectiveness of TransURL in complex multi-class classification tasks, indicating its potential as a promising solution for malicious URL detection in cybersecurity.
Figure 6: Detection results of baseline methods and TransURL on multiple classification dataset.
5.3 Cross-dataset Testing
@@
The results, as shown in Table 6, indicate a marked decline in performance of baseline methods on data not included in their training set, with URLNet’s accuracy dropping to 0.8823 and GramBeddings’ accuracy reducing to 0.5214 and 0.4647, respectively. In contrast, TransURL maintained high accuracy on external datasets, achieving an AUC of 0.9705. This underscores that the knowledge gained from one dataset by our method can be effectively generalized to others, even if the data was collected much later than TransURL’s pretraining period. This demonstrates the adaptability and long-term applicability of our approach to different data environments.
5.4 Evaluation against Adversarial Attacks
Figure 7: Area under ROC curve under adversarial attack.
@@ 5.4 Evaluation against Adversarial Attacks
As illustrated in Fig. 7 and Table 7, baseline methods experience a significant decline in performance under adversarial sample attacks. The accuracy of URLNet decreases to 76.10%, and that of GramBeddings to 70.18%, while our model maintains an accuracy above 90%, with an AUC of 98.39%, exceeding URLNet by about 20%. At a fixed FPR of 0.01, TransURL achieves a TPR of nearly 75%, more than double the TPR of baseline methods, both under 30%. These results indicate the robustness of our approach to adversarial sample attacks, suggesting increased effectiveness in preventing malicious attack evasion in real-world scenarios.
Table 7: Performance under adversarial attack.
@@
TransURL 0.9023 0.9738 0.8104 0.8846 0.9839
Table 8: Cross-dataset performance generalization.
@@
https://bafybeibfyqcvrjmwlpipqkdyt2xr46cea7ldciglcbybfwtk7cieugcj3e.
ipfs.infura-ipfs.io
@@
5.5 Case Study
We conduct a series of case studies applying our detection model to active malicious web pages to evaluate its practical utility. In November 2023, we crawled 30 active phishing URLs reported and verified on PhishTank and tested them with our model trained on 30% of the GramBeddings dataset. For comparison, tests were also conducted with the best-performing URLNet and GramBeddings models, trained on the same dataset. Results indicate that TransURL detected all malicious URLs with 100% accuracy, while GramBeddings misclassified 7 URLs, resulting in 76% accuracy, and URLNet misclassified 4 URLs, with an accuracy of 86%. Table 8 lists URLs that were incorrectly classified, in order to provide a detailed perspective on the performance of each model.
6 Discussion
@@
    End-to-End Architecture: TransURL is an end-to-end network utilizing pretrained CharBERT, which requires no manual feature initialization. It directly processes raw URLs and generates character-aware subword token embeddings. In contrast, previous studies typically necessitated manual initialization of character and word-level representations and relied on dual-path neural networks. Our approach streamlines the processing workflow while maintaining efficient feature extraction capabilities.
@@
    Evaluation Metrics: Our experiments show that TransURL significantly improves accuracy, robustness, and generalizability, consistently delivering stable and effective detection across various testing scenarios. Meanwhile, leading baseline methods, such as URLNet and Gramembedings, displayed clear weaknesses: URLNet struggled in class-imbalanced scenarios with small datasets, while Gramembedings faced significant performance fluctuation. Furthermore, these methods varied in their effectiveness in generalization and adversarial robustness tests. This highlights the need for a comprehensive performance evaluation system that goes beyond specific experimental setups.
@@
  4. 4.
    Computational Efficiency: A key consideration for the practical implementation of our proposed TransURL model is its computational efficiency, especially when deployed on resource-constrained devices such as endpoints. Given the intensive computational demands of Transformer-based models, the processing power required for real-time malicious URL detection can be substantial. To address this, we can employ optimization techniques such as model pruning and quantization, which significantly reduce memory usage without compromising detection accuracy. Additionally, we propose a hybrid approach where initial URL filtering is performed using a lightweight heuristic-based method. URLs flagged as potentially malicious are then subjected to more intensive scrutiny by the TransURL model. This layered approach ensures that the majority of URLs can be quickly processed with minimal computational overhead, while the TransURL model is reserved for cases where its advanced capabilities are most needed. These considerations make our approach viable for real-world applications, balancing the need for high detection accuracy with the constraints of endpoint devices. We will focus more on the solution to this problem in our future work.
7 Conclusion
We have proposed a novel transformer-based and pyramid feature learning system called TransURL for malicious URL detection. Our method effectively leverages knowledge transfer from pretrained models to URL contexts, dynamically integrates character and subword-level features, and incorporates three closely integrated feature learning modules for URL feature extraction. The key contributions of our approach are: 1) enabling end-to-end learning from raw URL strings without manual preprocessing; 2) adopting an interactive subword and character-level feature learning network architecture for improved character-aware subword representations; 3) conducting effective multi-level and multi-scale URL feature learning based on our proposed lightweight feature learning modules, addressing inherent limitations of the Transformer in local feature extraction and spatial awareness. We conduct extensive experiments on various URL datasets, demonstrating that our method consistently outperforms existing state-of-the-art baseline methods and produces stable decisions across scenarios. Furthermore, our method exhibits superior generalization and robustness in cross-dataset detection and adversarial sample attacks, enhancing its reliability in practical applications. We also provide a case study with comparative analysis to demonstrate the practical value of our method.
CRediT authorship contribution statement
Ruitong Liu: Conceptualization, Data curation, Formal analysis, Investigation. Yanbin Wang: Methodology, Writing– original draft, Writing– review & editing. Zhenhao Guo: Software, Validation, Conceptualization. Haitao Xu: Funding acquisition, Supervision. Wenrui Ma: Supervision. Fan Zhang: Project administration, Supervision.
Acknowledgements
The authors wish to express their sincere gratitude for the support received from the National Natural Science Foundation of China (NSFC) with the grant number 62272410.